OpenStudy (anonymous):

Suppose you deposit 2000 in a savings account that pays interest at an annual rate of 4%. If no money is added or withdrawn, determine how much with be in the account after 3 years. HELP!!! MEDAL

OpenStudy (anonymous):

4% = .04 right?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

This is a growth or decay equation but its complicated, do you know how?

OpenStudy (anonymous):

So it PAYS at a rate of 4% per YEAR for 3 YEARS

OpenStudy (anonymous):

1.04 is your growth per year and your time is 3 years

OpenStudy (anonymous):

Do you need me to explain how I got 1.04?

OpenStudy (anonymous):

Yes, if you can input the numbers into the equation? I want to see how you did it

OpenStudy (anonymous):

Yeah so first we know 4% is .04 If I want it to grow at a % rate I add a one to it, if I wouldnt do this then I would be left with 4% of my current number: 2000*.04 = 80 we know this wouldn't work because 80 < 2000. To be a growth rate we need a number > 2000. 2000* 1.04 = 2080 This is the growth for the first year. You could also do 2000*.04 + 2000 since 2000*.04 + 2000 = 2000*.04 + 2000 * 1 = 2000*1.04

OpenStudy (anonymous):

So in short to grow add a one to your percent

OpenStudy (anonymous):

So our first year total is 2080 Our second year would be 2080(1.04) and our third year would be (second year total)1.04

OpenStudy (anonymous):

AHHH omg

OpenStudy (anonymous):

So that's how! Great. THANK YOU SO MUCH!!!!!! I can do my homework with ease. Thank you :) <3

OpenStudy (anonymous):

Yeah no problem

OpenStudy (ahvei):

ok here is my answer.... So our first year total is 2080 Our second year would be 2080 and our third year would be 1.04

OpenStudy (ahvei):

hope this helps

OpenStudy (anonymous):

Be nice guys

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

Thanks @swagmaster47 and please behave @ahvei

OpenStudy (ahvei):

ok fine sorry

OpenStudy (anonymous):

Thanks :), have a great day!